# Vec2Face: Scaling Face Dataset Generation with Loosely Constrained Vectors

### v1

@@ 3.2 Sampling identities and face images
where $v_{im_{k}}$ is the $k$th vector for an identity, $k=1,...,m$. $\mathcal{N}(0,\sigma)$ is a Gaussian distribution with 00 mean and $\sigma$ variance. To ensure the identity-consistency after perturbation, the similarity values between perturbed vectors and their identity vectors are $\geq 0.5$.
Technically speaking, the above steps allow us to synthesize an unlimited number of identities and images using loosely constrained vectors. Hyperparameter $\tau$ is selected such that identities are well-separated, i.e., high inter-class variation, although it does not act as a strong filter in practice; the selection of hyperparameter $\sigma$ should enable the generated images for an identity to have large enough variance while staying on the same identity, i.e., high intra-class variation.
3.3 Explicit face attribute control
By default, Vec2Face generates images without explicit attribute control. This might lead to some attributes being under-represented, e.g., too few profile head poses. We introduce attribute operation, or AttrOP, to guide vector perturbation so that additional images of desired attributes can be generated. This work focuses on face pose and image quality.
Inspired by (Singh et al., 2023), AttrOP controls the attributes of the generated images by simply adjusting the values in the feature vectors via gradient descent, a process shown in Algorithm 1. Specifically, we first set a target image quality $Q$ and pose angle $P$ and prepare a pretrained pose evaluation model $M_{pose}$ and a pretrained quality evaluation model $M_{quality}$. Then, given an identity vector $v_{id}$ and a perturbed identity vector $v_{im}$, we generate a face image from an adjusted vector $v^{\prime}_{im}$ and compute the following loss functions:
@@ 3.4 Discussion
Novelty statement. Generating large-scale training sets presents high demands on inter-class separability and intra-class variation. Controlling the identity by either a coefficient or hard class labels is not effective. Conditioning on identity vectors or clustering generated images with a FR model could effectively generate well-separated identities but needs an additional model to address the shortage of intra-class variation, which is highly dependent on the reference images. This approach lacks flexibility in multi-attribute control. The face features contain all the information of faces, so Vec2Face is designed to learn face generation purely from the given face features. The advantage is that, once the model learns how to draw the target image with the given feature, the output can be easily controlled to add desired information to the vector (i.e., adjusting the values).
Why can Vec2Face generate images like in Fig. 1? The FR model maps images to a hyperspherical space and clusters them. This results in, high similarity for image features from the same identity and low similarity for image features from different identities. Vec2Face learns how to generate face images based on the information in the vectors. Increasing the degree of the perturbation added to the vector gradually changes the vector direction, which results in faces from different identities.
How can Vec2Face increase intra-class separability and inter-class variations? A face feature vector has both identity and attribute information. Hence, controlling the cosine similarity between sampled vectors can enforce separate identities and adding small perturbations to sampled vectors could change the attributes while maintaining the identity. Moreover, AttrOP allows to more easily get the vectors to generate images with desired attributes.
Can other structures work? A diffusion model conditioned on image features may achieve similar function. However, the results in Appendix  A.1 show that the initial noise image may exert a stronger influence on the final output than the conditioned image/identity features, which fails to preserve the identity. Our structure is image-feature-only, and the image feature is used to replace the masked-out rows in the fMAE, ensuring less noise from other factors. Thus, unless there are other special designs, our image-feature-only structure is likely to be more appropriate for this task.
What is an identity? An interesting philosophical question would be how to define an identity. Our research suggests that generated identities can be continuous, where a threshold $\tau$ is used to determine the boundary between different identities. It will be interesting to study how this computational definition of identities connects with and differs from philosophical understanding in future.
@@ 4.1 Datasets and evaluation protocol
Vec2Face training set. The training data consists of 1M images and their features from 50K randomly sampled identities in the WebFace4M (Zhu et al., 2023) dataset, where 20 images per identity provides sufficient data for Vec2Face to learn the pattern.
FR model training set. The FR model for feature extraction is ArcFace-R100 trained with Glint360K (An et al., 2021) dataset. The training configurations are the same as used in the insightFace repository [^1] https://github.com/deepinsight/insightface/.
Real-world test sets. The FR models are evaluated by five commonly used test sets. LFW (Huang et al., 2008) tests the FR model in a general case. CFP-FP (Sengupta et al., 2016) and CPLFW (Zheng & Deng, 2018) test the FR model on pose variation. AgeDB (Moschoglou et al., 2017) and CALFW (Zheng et al., 2017) challenge the FR model with large age gap. Besides these five, Hadrian (Wu et al., 2024a) and Eclipse (Wu et al., 2024a) are used for intra-class variation test, where Hadrian pairs emphasize facial hair difference and Eclipse pairs differ on face exposure. The images in both datasets are indoor and high quality originating from MORPH (Ricanek & Tesafaye, 2006). Ability to discriminate between similar-looking persons is tested by two datasets designed for that purpose (Deng et al., 2017; Thom et al., 2023).
Evaluation protocol. The image pairs in each test set are split into 10 folds. The average value of the 10-fold cross-validation is used to evaluate the model performance, as suggested in (Huang et al., 2008).
@@ 4.2 Experiment details
Vec2Face training: We used the default ViT-Base as the backbone to form the $MAE_{f}$. Since the image size of the commonly used FR training sets is 112x112 and the image decoder is a four-layer, double-sized architecture, the feature map size is set to 49x768. After stacking the projected image features, the entire feature map becomes 50x768. The optimizer is AdamW (Loshchilov & Hutter, 2017), with a learning rate of 4e-5 and a batch size of 32 per GPU. We use approximately 10 RTX6000 GPUs for each training run. The patch-based discriminator is enabled after 1000 epochs.
Dataset generation. We sample separable identity vectors by PCA learned on identity vectors of MS1MV2 (Deng et al., 2019). We perturb the vectors with the noise sampled from three Gaussian distributions. Specifically, 40% from $\mathcal{N}(0,0.3)$, 40% from $\mathcal{N}(0,0.5)$, and 20% from $\mathcal{N}(0,0.7)$ for each vector. For extreme pose image generation, we generate 30 images for each identity and randomly replace the existing ones. Specifically, the target pose $\mathcal{P}$ of 20 images is 60 and 10 images is 85. Meanwhile, we observe an image quality degradation with pure pose control, so we control the image quality with a target value 27. The iterations $T$ is initialized as 5, but we add more conditions to the vanilla AttrOP algorithm to ensure the images have the designated attributes.
AttrOP details. MagFace-R100 (Meng et al., 2021) is used as the quality evaluation model, SixDRepNet (Hempel et al., 2024) is used as the pose evaluation model, and ArcFace-R100 is used as the identity evaluation model.
FR dataset evaluation metric. The quality of the synthetic datasets is evaluated by the performance of the trained model on five test sets. The standard training dataset size is 0.5M images from 10K identities. Unless otherwise specified, the backbone is SE-IR50 (He et al., 2016), the recognition loss is ArcFace (Deng et al., 2019), other configuration details are in the Appendix.
@@ 4.3 Main Evaluation
Training sets # images LFW CFP-FP CPLFW AgeDB CALFW Avg.
IDiff-FaceBoutros et al. (2023a) 0.5M 98.00 85.47 80.45 86.43 90.65 88.20
DCFaceKim et al. (2023) 0.5M 98.55 85.33 82.62 89.70 91.60 89.56
Arc2FacePapantoniou et al. (2024) 0.5M 98.81 91.87 85.16 90.18 92.63 91.73
DigiFaceBae et al. (2023) 1M 95.40 87.40 78.87 76.97 78.62 83.45
SynFaceQiu et al. (2021) 0.5M 91.93 75.03 70.43 61.63 74.73 74.75
SFaceBoutros et al. (2022a) 0.6M 91.87 73.86 73.20 71.68 77.93 77.71
IDnetKolf et al. (2023) 0.5M 92.58 75.40 74.25 63.88 79.90 79.13
ExFaceGANBoutros et al. (2023b) 0.5M 93.50 73.84 71.60 78.92 82.98 80.17
SFace2Boutros et al. (2024) 0.6M 95.60 77.11 74.60 77.37 83.40 81.62
Langevin-DiscoGeissbühler et al. (2024) 0.6M 96.60 73.89 74.77 80.70 87.77 82.75
HSFace10K (Ours) 0.5M 98.87 88.97 85.47 93.12 93.57 92.00
CASIA-WebFace (Real) 0.49M 99.38 96.91 89.78 94.50 93.35 94.79
Table 1: Comparison of existing synthetic datasets on five real-world test sets. $\dagger$, $\star$, and $\diamond$ represent diffusion, 3D rendering, and GAN approaches, respectively, for constructing these datasets. We also list the results of training on a real-world dataset CASIA-WebFace.
Comparison with state-of-the-art synthetic datasets at the scale of 0.5M-0.6M images. In Table 1, we compare FR accuracy of models trained with datasets synthesized by different methods such as diffusion models, GANs, and 3D rendering. We have the following observations. First, HSFace10K synthesized by Vec2Face yields very competitive accuracy: 98.87%, 88.97%, 85.47%, 93.12%, 93.57% on LFW, CFP-FP, CPLFW, AgeDB, and CALFW, respectively, and 92.00% on average. Our method is only lower than Arc2Face on the CFP-FP dataset (88.97% vs. 91.87%) and is the state of the art on all the other datasets and average. Second, on the CALFW dataset, the accuracy of HSFace10K is higher than, or at least competitive with that of CASIA-WebFace (93.57% vs. 93.32%). To our knowledge, this is the first time that a model trained with synthetic data outperforms those trained with same-size real data. Third, generally speaking, GAN-based methods, while being prevalent, are not as competitive as diffusion-based and 3D rendering methods. Our method is GAN-based but has very strong performance.
Methods # images LFW CFP-FP CPLFW AgeDB CALFW Avg.
@@
Table 2: Impact of scaling the proposed HSFace dataset to 1M images (20K IDs), 5M images (100K IDs), 10M images (200K IDs), 15M images (300K IDs). Continued improvement is observed. We also list the performance obtained by training on the real-world dataset CASIA-WebFace and its combination with HSFace10K. The latter combination yields even higher accuracy.
Effectiveness of scaling up the proposed HSFace dataset. Vec2Face can easily generate a large number of distinct identity images by sampling more identity vectors which have a cosine similarity $\leq\tau$ with any other identity vectors. Specifically, we generate 10k, 20k, 100k, 200k and 300k IDs, where each identity has 50 images. Results of models trained on these datasets are shown in Table 2.
It is clear that scaling up our synthetic dataset leads to consistent accuracy improvements: from 92.00%, to 92.47%, 92.97%, 93.22%, and 93.52%. The largest synthetic face dataset prior to this paper has 60k identities and 1.2M images (Papantoniou et al., 2024). Our dataset reaches 12.5 times larger while still giving steady improvement. These results clearly demonstrate the advantage of Vec2Face in data scaling. Also note that Vec2Face is only trained with 50K real-world identities. We speculate that a Vec2Face model trained with larger initial real data could bring about further improvements.
Merging synthetic and real-world training sets. In Table 2, when we merge HSFace10K and CASIA-WebFace, we observe improvement over using either dataset alone for training. We obtain 95.50% average accuracy, which is higher than 92.00% (HSFace10K alone) and 94.79% (CASIA-WebFace alone). This indicates that HSFace10k has good quality and is complementary to real data.
Effectiveness of attribute control. Images generated by Vec2Face are mostly near-frontal, which can be attributed to such images being most frequent in WebFace4M, the dataset used to train Vec2Face. Using AttrOP, we synthesize faces with some profile poses where we use target Yaw angles 60 and 85. Results are shown in Table 4. We observe that adding faces with 60 poses leads to 3.46% improvement, while adding 85 gives further 0.5% improvement. These results demonstrate the effectiveness of attribute control. Note that this is achieved with Vec2Face only, instead of resorting to external models as in previous works (Kim et al., 2023; Papantoniou et al., 2024).
Models Computing cost FID
@@
Table 3: Computing cost and FID measurement of Arc2Face (Papantoniou et al., 2024) and Vec2Face.
Image generation cost and quality of reconstruction. The resource requirement for image generation is an important metric, especially in this large model era. We compare our method with the state-of-the-art face generation model. Experimental details: 1) both models are tested on a single Titan-Xp, 2) the batch size is 8, 3) a 4-steps sampler is used for Arc2Face. Table 3 shows that our method is 311x faster than Arc2Face. To test the quality of generated images, we use FID to measure the distance between the original LFW and Hadrian images, and the reconstructed images from Arc2Face and Vec2Face. The images are measured at 112x112 resolution. Our model slightly outperforms the Arc2Face results, in size of model, speed of generation, and fidelity to original images, as shown in Table 3. The similarity measurement and examples of reconstructed images are provided in the Appendix A.1.
Attribute control LFW CFP-FP CPLFW AgeDB CALFW Avg.
- 98.27 76.56 81.70 90.75 92.92 88.04
+ face quality 27 98.55 83.27 83.68 91.12 93.27 89.98
+ Yaw angle 60 98.62 86.46 85.75 92.85 93.80 91.50
+ Yaw angle 85 98.87 88.97 85.47 93.12 93.57 92.00
Table 4: Impact of face quality and pose control via AttrOP on FR performance. Each of the training sets has 0.5M images. The accuracy improvement is observed when more controls are added for image quality and pose angles, especially on pose-oriented test sets, CFP-FP and CPLFW.

### v4

@@ 3.2 Inference: Sampling vectors to generate identities and their face images
where $v_{im_{k}}$ is the $k$th perturbed vector for an identity vector $v_{n}$. $\mathcal{N}(0,\sigma)$ is a Gaussian distribution with 00 mean and $\sigma$ variance. To ensure the identity-consistency after perturbation, the similarity values between perturbed vectors and their identity vectors are $\geq 0.5$.
Technically, the above steps allow us to synthesize an unlimited number of identities and their images using loosely constrained vectors via Vec2Face. Hyperparameter $\tau$ is selected such that identities are well-separated, i.e., high inter-class variation, although it does not act as a strong filter in practice. The selection of hyperparameter $\sigma$ should enable the generated images for an identity to have large enough variance while staying on the same identity, i.e., high intra-class variation.
3.3 Inference: Explicit face attribute control
By default, Vec2Face generates images without explicit attribute control. This might lead to some attributes being under-represented, e.g., too few profile head poses. To address this, we introduce attribute operation, or AttrOP, to guide vector perturbation so that additional images of desired attributes can be generated. This work focuses on face pose and image quality.
Inspired by (Singh et al., 2023), AttrOP controls the attributes of the generated images by simply adjusting the values in the feature vectors via gradient descent, a process shown in Algorithm 1. Specifically, we first set a target image quality $Q$ and pose angle $P$ and prepare a pretrained pose evaluation model $M_{pose}$, a pretrained quality evaluation model $M_{quality}$, and a FR model $M_{FR}$. Then, given an identity vector $v_{id}$ and a perturbed identity vector $v_{im}$, we generate a face image from an adjusted vector $v^{\prime}_{im}$ and compute the following loss functions:
@@ 3.4 Discussion
Novelty statement. Vec2Face is novel in three aspects. First, it is new to add small/large perturbations to input vectors to generate images of the same/different IDs. In comparison, existing works typically use fixed ID vectors without an explicit mechanism to separate different IDs. Second, as a minor contribution, our GAN architecture realizes the above function, where the fMAE component prevents overfitting during training, similar to VAE (Kingma & Welling, 2013). Third, Vec2Face can seamlessly support attribute control during inference, while existing methods have to rely on external models for this purpose.
Why can Vec2Face generate images like in Fig. 1? The FR model maps images to a hyperspherical space, where features of the same identity are clustered, and those of different identities are farther apart (Deng et al., 2019). Using such features as input and supervised by the original images, Vec2Face learns how to generate face images based on characteristics of these vectors. As such, increasing the degree of the perturbation added to the vector gradually changes the vector direction, which results in faces of the same and then different identities.
How can Vec2Face increase inter-class separability and intra-class variations? A face feature vector reflects both identity and attributes. Hence, controlling the cosine similarity between sampled vectors can enforce separate identities, and adding small perturbations to sampled vectors could change the attributes while maintaining the identity. Moreover, AttrOP allows us to easily obtain useful vectors to generate images with desired attributes, further increasing intra-class variations.
Can other structures work? A diffusion model conditioned on image features may achieve similar function. However, the results in Appendix A.1 show that the initial noise image may exert a stronger influence on the final output than the conditioned image/identity features. This compromises identity-preserving capability. In comparison, the stochasticity of our method mainly comes from the image feature. fMAE has some randomness in testing, but has a very minor effect on generated images. That said, it would be interesting to design more effective architectures in the future.
What is an identity? An interesting philosophical question would be how to define an identity. Our research suggests that generated identities can be continuous, where a threshold $\tau$ is used to determine the boundary between different identities. We conducted a simple experiment for this in Section 4.4. It will be interesting to deeply study how this computational definition of identities connects with and differs from philosophical understanding in the future.
@@ 4.1 Datasets and evaluation protocol
Vec2Face training set. The training data consists of 1M images and their features from 50K randomly sampled identities in WebFace4M (Zhu et al., 2023), where the images features are extracted by an ArcFace-R100 model pretrained on Glint360K (An et al., 2021). Unless otherwise specified, this model is used for feature extraction throughout this paper.
Real-world test sets. There are eleven test sets used to compare the synthetic and real datasets on FR model training. LFW (Huang et al., 2008) tests the FR model in a general case. CFP-FP (Sengupta et al., 2016) and CPLFW (Zheng & Deng, 2018) test the FR model on pose variation. AgeDB (Moschoglou et al., 2017) and CALFW (Zheng et al., 2017) challenge the FR model with large age gap. These five test sets are used for a general accuracy comparison across synthetic datasets and a real dataset. Besides these five, Hadrian (Wu et al., 2024a) and Eclipse (Wu et al., 2024a) are used for intra-class variation evaluation, where Hadrian pairs emphasize facial hair difference and Eclipse pairs differ on face exposure. The images in both datasets are indoor and high quality originating from MORPH (Ricanek & Tesafaye, 2006). In addition, SLLFW (Deng et al., 2017) and DoppelVer (Thom et al., 2023), are used to evaluate the identity definition used in the existing works (including ours). These two test sets have similar-looking pairs. Lastly, IJBB (Whitelam et al., 2017) and IJBC (Maze et al., 2018) are more challenging test sets that are closer to the real-world scenario.
Evaluation protocol of test sets. Besides IJBB and IJBC, the aforementioned test sets use the same evaluation protocol suggested in (Huang et al., 2008). Following the previous work (Deng et al., 2019), we use TPR@FPR=1e-4 to measure the model performance on IJBB and IJBC.
@@ 4.2 Experiment details
Vec2Face model training: We used the default ViT-Base as the backbone to form the fMAE. Since the image size of the commonly used FR training sets is 112x112 and the image decoder is a four-layer, double-sized architecture, the feature map size is set to 49x768, so that this feature map can be reshaped to 7x7x768 for image decoder to decode a 112x112x3 image. The optimizer is AdamW (Loshchilov & Hutter, 2017), with a learning rate of 4e-5 and a batch size of 32 per GPU. We use approximately 10 RTX6000 GPUs for each training run.
Synthetic dataset generation. We obtain well-separated identity vectors by PCA learned on face feature vectors of MS1MV2 (Deng et al., 2019), details are in Section 3.2. To get vectors for image generation, we perturb the identity vectors with the noise sampled from three Gaussian distributions. Specifically, 50 perturbed vectors are sampled for each identity, where 40% from $\mathcal{N}(0,0.3)$, 40% from $\mathcal{N}(0,0.5)$, and 20% from $\mathcal{N}(0,0.7)$. But this resulting dataset suffers from the low pose variation, which causes bad performance on CFP-FP and CPLFW. To increase pose variation, we generate 30 images with large yaw angle via AttrOP for each identity and randomly replace the existing ones. In detail, the target pose $\mathcal{P}$ of 20 images is 60 and 10 images is 85. Meanwhile, we add the image quality control, $Q=27$, to mitigate the quality degradation during pose adjustment.
AttrOP details. The MagFace-R100 (Meng et al., 2021), SixDRepNet (Hempel et al., 2024), and ArcFace-R100 are used as $M_{quality}$, $M_{pose}$, and $M_{FR}$ in Algorithm 1. $T$ is set to 5.
Face synthesizing method comparison settings. The standard training dataset size is 0.5M images from 10K identities. Unless otherwise specified, the backbone is SE-IR50 (He et al., 2016), the recognition loss is ArcFace (Deng et al., 2019), and other configuration details are in the Appendix A.9.
@@ 4.3 Main Evaluation
Training sets # ims LFW CFP-FP CPLFW AgeDB CALFW Avg. IJBB IJBC
IDiff-Face (Boutros et al., 2023a) 0.5M 98.00 85.47 80.45 86.43 90.65 88.20 62.13 63.61
DCFace (Kim et al., 2023) 0.5M 98.55 85.33 82.62 89.70 91.60 89.56 66.47 69.92
ID3 (Li et al., 2024a) 0.5M 97.68 86.84 82.77 91.00 90.73 89.80 - -
Arc2Face (Papantoniou et al., 2024) 0.5M 98.81 91.87 85.16 90.18 92.63 91.73 - -
DigiFace (Bae et al., 2023) 1M 95.40 87.40 78.87 76.97 78.62 83.45 5.42 6.10
SynFace (Qiu et al., 2021) 0.5M 91.93 75.03 70.43 61.63 74.73 74.75 - -
SFace (Boutros et al., 2022a) 0.6M 91.87 73.86 73.20 71.68 77.93 77.71 37.90 41.94
IDnet (Kolf et al., 2023) 0.5M 92.58 75.40 74.25 63.88 79.90 79.13 - -
ExFaceGAN (Boutros et al., 2023b) 0.5M 93.50 73.84 71.60 78.92 82.98 80.17 - -
SFace2 (Boutros et al., 2024) 0.6M 95.60 77.11 74.60 77.37 83.40 81.62 - -
HSFace10K (Ours) 0.5M 98.87 88.97 85.47 93.12 93.57 92.00 83.82 86.96
CASIA-WebFace (Real) 0.49M 99.38 96.91 89.78 94.50 93.35 94.79 78.71 83.44
Table 1: Comparison of existing synthetic datasets on seven real-world test sets. $\dagger$, $\star$, and $\diamond$ represent diffusion, 3D rendering, and GAN approaches, respectively, for constructing these datasets. We also list the results of training on a real-world dataset CASIA-WebFace. Since previous works do not provide the IJB results, we re-train the FR models and report the accuracy on available datasets.
Comparison with state-of-the-art synthetic datasets at the scale of 0.5M-0.6M images. In Table 1, we compare FR accuracy of models trained with datasets synthesized by different methods, i.e., diffusion models, GANs, and 3D rendering. We have the following observations. First, HSFace10K synthesized by Vec2Face yields very competitive accuracy: 98.87%, 88.97%, 85.47%, 93.12%, 93.57% on LFW, CFP-FP, CPLFW, AgeDB, and CALFW, respectively, and 92.00% on average. Our method is only lower than Arc2Face on the CFP-FP dataset (88.97% vs. 91.87%) and is the state of the art on all the other datasets and average. Second, on the CALFW, IJBB, and IJBC datasets, the accuracy of HSFace10K is higher than that of CASIA-WebFace (Yi et al., 2014). To our knowledge, this is the first time that a model trained with synthetic data outperforms that trained with same-scale real data. Third, generally speaking, GAN-based methods, while being prevalent, are not as competitive as diffusion-based and 3D rendering methods. Our method is GAN-based but outperforms the other methods.
Datasets # images LFW CFP-FP CPLFW AgeDB CALFW Avg.
@@
Table 2: Impact of scaling the proposed HSFace dataset to 1M images (20K IDs), 5M images (100K IDs), 10M images (200K IDs), 15M images (300K IDs), 20M images (400K IDs). Continued improvement is observed until 300K IDs. We also list the performance obtained by training on the real-world dataset CASIA-WebFace and its combination with HSFace10K. The latter combination yields even higher accuracy.
Effectiveness of scaling up the proposed HSFace dataset. Vec2Face can easily generate a large-scale dataset by sampling more identity vectors which have a cosine similarity $\leq\tau$ with any other identity vectors. To test the efficacy of Vec2Face on scaling, we generate six datasets with increased identity numbers from 10K to 400K, where each identity has 50 images. Results of models trained on these datasets are shown in Table 2.
It is clear that scaling up our synthetic dataset leads to consistent accuracy improvements until 300K IDs: from 92.00%, to 92.47%, 92.97%, 93.22%, 93.52%. Notably, our HSFace300K is 12.5 times larger than the largest synthetic dataset prior to this paper while still giving steady accuracy improvement. These results clearly demonstrate the advantage of Vec2Face in data scaling. Also note that Vec2Face is only trained with 1M real-world images. We speculate that a Vec2Face model trained with larger initial real data could bring further improvements.
Merging synthetic and real-world training sets. In Table 2, after merging HSFace10K and CASIA-WebFace, we observe improvement over using either dataset alone for training. We obtain 95.50% average accuracy, which is higher than 92.00% (HSFace10K alone) and 94.79% (CASIA-WebFace alone). This indicates that HSFace10K has good quality and is complementary to real data. Additional experiments are in the Appendix A.9.
Attr. control LFW CFP-FP CPLFW AgeDB CALFW Avg.
- 98.27 76.56 81.70 90.75 92.92 88.04
+ Quality 27 98.55 83.27 83.68 91.12 93.27 89.98
+ Angle 60 98.62 86.46 85.75 92.85 93.80 91.50
+ Angle 85 98.87 88.97 85.47 93.12 93.57 92.00
Table 3: Impact of face quality and pose control via AttrOP in FR accuracy. Each of the training sets has 0.5M images. The accuracy improvement is observed when more controls are added for image quality and pose angles, especially on pose-oriented test sets, CFP-FP and CPLFW.
Effectiveness of attribute control. Images generated by Vec2Face are mostly near-frontal, which can be attributed to such images being most frequent in the training set. Using AttrOP, we synthesize faces with some profile poses where we use target Yaw angles 60 and 85. Results are shown in Table 3. We observe that adding faces with 60 poses leads to 3.46% improvement, while adding 85 gives further 0.5% improvement. These results demonstrate the effectiveness of attribute control. Unlike prior works (Kim et al., 2023; Papantoniou et al., 2024), this is achieved by a single model.
Methods Computing cost FID
@@
Table 4: Computing cost and FID measurement of Arc2Face and Vec2Face.
Image generation cost and quality of reconstruction. The resource requirement for image generation is an important metric, especially in this large model era. We compare our method with a state-of-the-art face generation model. Experimental details: 1) both models are tested on a single Titan-Xp, 2) the batch size is 8, 3) a 4-step scheduler (Luo et al., 2023) is used for Arc2Face. Table 4 shows that our method is 311x faster than Arc2Face. We also use FID to measure the distance between the original LFW and Hadrian images, and the reconstructed images from Arc2Face and Vec2Face, where the images are measured at 112x112 resolution. Our model slightly outperforms the Arc2Face results, in size of model, speed of generation, and fidelity to original images , as shown in Table 4. More results are in Appendix A.1.